# Sample Compression Unleashed: New Generalization Bounds for Real Valued Losses

### v2

@@ A.1.3 Regression problems
We trained decision trees and forests on the datasets, using P2L to train the forests on one datapoint at a time. We trained the models until their validation loss hasn’t decreased for 10 or 20 epochs. We summarize this idea in Algorithm 2. We denote the RMSE as $\ell^{\textrm{RMSE}}(h,\boldsymbol{x},y)=\sqrt{(h(\boldsymbol{x})-y)^{2}}$ and the empirical risk on the dataset
@@
Input : $T$, the number of iterations before stopping.
Initialize : $S_{\operatorname{\mathbf{i}}}=\emptyset$.
Initialize : $h_{\operatorname{\mathbf{i}}}=h_{0}$.
Initialize : $\operatorname{\mathcal{L}}_{\textrm{BEST}}=\infty.$
Initialize : $\texttt{COUNTER}=0$.
Initialize : $(\overline{\boldsymbol{x}},\overline{y})=\operatorname*{arg\,max}_{(\boldsymbol{x},y)\in S}\ell^{\textrm{RMSE}}(h_{0},\boldsymbol{x},y)$
@@
  • Maximum depth of the trees : $[5,10]$
  • Minimum samples to split : $[2,3,4]$
  • Minimum samples to create a leaf : $[1,2,3]$
  • Cost-Complexity pruning parameter : $[0.0,0.05,0.1,0.2,0.5,1,2]$
  • Number of epochs before stopping : $[10,20]$
@@
For the decision forests, we choose the number of estimators in $[50,100]$. For the baselines, we train the same model with the same hyperparameters on the whole dataset.
The results for the decision trees can be found in Table 11. Using only P2L to train the trees leads to underfitted trees, as the model is not complex enough to use only a few datapoints to train a complete model.
When computing the bounds, we need to bound the loss, as the RMSE is not bounded. However, the regression trees cannot predict a value bigger (respectively smaller) than the biggest (respectively smallest) target value found in the dataset. Thus, the loss is bounded by the biggest and smallest target values found in the dataset $S$. To compute the $\mathrm{kl}$ bound, we add an assumption on the data-generating distribution $\operatorname{\mathcal{D}}$, which is that the target values of $\operatorname{\mathcal{D}}$ are bounded by :
$\min_{(\boldsymbol{x},y)\sim S}y-p\quantity(\max_{(\boldsymbol{x},y)\sim S}y-\min_{(\boldsymbol{x},y)\sim S}y)\leq\min_{(\boldsymbol{x},y)\sim\operatorname{\mathcal{D}}}y\leq\max_{(\boldsymbol{x},y)\sim\operatorname{\mathcal{D}}}y\leq\max_{(\boldsymbol{x},y)\sim S}y+p\quantity(\max_{(\boldsymbol{x},y)\sim S}y-\min_{(\boldsymbol{x},y)\sim S}y)$
where $p\in[0\%,100\%]$. For the experiments, we choose $p=10\%$. If $\min_{(\boldsymbol{x},y)\sim S}y=0$, then we simply lower bound by 0.
To compute the linear bound, we assume that the distribution is $\varsigma^{2}$-sub-Gaussian with $\varsigma=\frac{1}{2}\quantity(\max_{(\boldsymbol{x},y)\sim S}y-\min_{(\boldsymbol{x},y)\sim S}y)$.
These assumptions are restrictive and we would rather have no assumption on the data-generating distribution. In some settings, we can remove the assumption on the dataset by having expert knowledge on the distribution. For example, if you predict a probability, the target domain is $[0,1]$.
We report the lower bounds, minimum values in the training set, maximum values in the training set and upper bounds for each dataset in Table 12.
@@
Table 11: Results for the decision trees trained using P2L. We report the RMSE achieved by the models and the generalization bounds on the RMSE.
@@
Dataset Train loss Validation loss Test loss $\mathrm{kl}$ bound Linear bound $\operatorname{|\mathbf{i}|}$ Baseline test loss $\ell^{\max}$
Powerplant 11.66$\pm$3.00 11.83$\pm$3.12 12.00$\pm$3.04 23.40±1.59 24.15±1.54 72.80$\pm$38.32 4.07$\pm$0.13 90.6
Infrared 0.77±0.11 0.80±0.08 0.76±0.13 1.63±0.14 1.55±0.11 12.80±0.40 0.27$\pm$0.03 4.26
Airfoil 11.02$\pm$1.71 10.89$\pm$1.38 11.10$\pm$1.93 18.87±1.97 18.14±1.75 12.20$\pm$0.40 3.01$\pm$0.19 45.13
Parkinson 14.75$\pm$1.86 14.53$\pm$2.19 14.90$\pm$2.22 18.86±1.95 18.28±1.86 12.00$\pm$0.00 3.20$\pm$0.15 41.37
Concrete 26.44$\pm$1.99 27.40$\pm$1.85 27.08$\pm$1.68 46.56±1.74 45.00±1.72 14.00$\pm$1.26 6.22$\pm$0.91 90.63
@@
Table 12: Minimum and maximum values used to compute the bound on regression problems.
Dataset Lower bound Minimum value Maximum value Upper bound
@@ A.1.4 Amazon Polarity
We trained DistilBERT on the Amazon Reviews Polarity dataset. We use a subset of 10% of the real dataset, amounting to 360000 datapoints. 180000 are used for pretraining the model, 144000 for training and 36000 for validation. We use the given test set of 400000 datapoints. Using P2L, we add 32 datapoints at a time to the compression set and stop the training when the validation loss hasn’t decreased in 20 iterations. For both P2L and the baseline, which was trained for 200 epochs or until it reached 0 errors on the training dataset, we use the following hyperparameter grid :
  • Number of pretraining epochs : $[2,5]$
  • Pretraining learning rate : 2e-5
  • Dropout probability : $[0.1,0.2]$
  • Training learning rate : [1e-6, 1e-7, 1e-8]
For this model, we trained using only one seed.
@@ Appendix B Theoretical results from the literature
$\forall\operatorname{\mathbf{i}}\in I,\forall\sigma\in M(\operatorname{\mathbf{i}})\,:$ $\ R_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma))\leq 1-\exp\quantity(\frac{-1}{n-\operatorname{|\mathbf{i}|}-\kappa}\quantity[\ln\matrixquantity(n-\operatorname{|\mathbf{i}|}\\ \kappa)+\ln\quantity(\frac{\matrixquantity(n\\ \operatorname{|\mathbf{i}|})}{\zeta(\operatorname{|\mathbf{i}|})P_{M(\operatorname{\mathbf{i}})}(\sigma)\delta})])\,,$
with $\kappa=nR_{S_{\operatorname{\mathbf{i}}^{c}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma))$.
@@ C.1 Proof of the main result
See 3 Before proving Theorem 3, we need to restate Chernoff’s bound in a way that will be useful to prove Theorem 3.
@@
$\operatorname*{\mathbb{P}}\quantity(X\leq\frac{1}{t}\quantity[\ln\operatorname*{\mathbb{E}}e^{tX}+\ln\frac{1}{\delta}])\geq 1-\delta.$
@@
Chernoff’s bound states that for a random variable $X$, any $t>0$ and $\epsilon>0$, we have :
$\operatorname*{\mathbb{P}}\quantity(X>\epsilon)\leq e^{-t\epsilon}\operatorname*{\mathbb{E}}e^{tX}$
By choosing $\delta=e^{-t\epsilon}\operatorname*{\mathbb{E}}e^{tX}$, we have :
@@
$\iff$ $\epsilon=\frac{1}{t}\quantity[\ln\operatorname*{\mathbb{E}}e^{tX}+\ln\frac{1}{\delta}]$
Thus, we have :
@@
We start by defining a sample-compressed set. Given a dataset $S\sim\operatorname{\mathcal{D}}^{n}$ and $\operatorname{\mathcal{H}}$ a predictor set, we consider the following subset of $\operatorname{\mathcal{H}}$, that contains only sample-compressed predictors :
^$\widehat{\operatorname{\mathcal{H}}}_{S}\coloneqq\left\{\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma)|\operatorname{\mathbf{i}}\in I,\sigma\in M(\operatorname{\mathbf{i}})\right\}\subseteq\operatorname{\mathcal{H}}.$
For any vector of indices $\operatorname{\mathbf{i}}\in I$ and a message $\sigma\in M(\operatorname{\mathbf{i}})$, when given a dataset $S$, we fully define a predictor $\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma)\in\widehat{\operatorname{\mathcal{H}}}_{S}$. For a specific pair $(\operatorname{\mathbf{i}},\sigma)$, let’s study the value of $\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{S_{\operatorname{\mathbf{i}}^{c}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma)))$, a realization of a random variable of mean
@@
With $\delta_{\operatorname{\mathbf{i}}}^{\sigma}\in(0,1)$ and $t>0$, using Chernoff’s bound as stated in Lemma 11, we have :
@@
$\geq 1-\delta_{\operatorname{\mathbf{i}}}^{\sigma}$
Using the union bound, we get a bound that is valid for all pairs $(\operatorname{\mathbf{i}},\sigma)$ simultaneously,
$\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\forall\operatorname{\mathbf{i}}\in I,\sigma\in M(\operatorname{\mathbf{i}}):\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{S_{\operatorname{\mathbf{i}}^{c}}}(h_{\operatorname{\mathbf{i}}}^{\sigma}),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(h_{\operatorname{\mathbf{i}}}^{\sigma}))\leq\frac{1}{t}\quantity[\ln\operatorname*{\mathbb{E}}_{T_{\operatorname{\mathbf{i}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{i}|}}}\operatorname*{\mathbb{E}}_{T_{\operatorname{\mathbf{i}}^{c}}\sim\operatorname{\mathcal{D}}^{n-\operatorname{|\mathbf{i}|}}}e^{t\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{T_{\operatorname{\mathbf{i}}^{c}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{i}}},\sigma)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{i}}},\sigma)))}+\ln\frac{1}{\delta_{\operatorname{\mathbf{i}}}^{\sigma}}])$
$\geq 1-\sum_{\operatorname{\mathbf{i}}\in I}\sum_{\sigma\in M(\operatorname{\mathbf{i}})}\delta_{\operatorname{\mathbf{i}}}^{\sigma}.$ (9)
To obtain a valid bound, we will need to either compute or bound the following term :
@@
As the set $T$ is a realization of a $n-\operatorname{|\mathbf{i}|}$ datapoints, choosing $t=n-\operatorname{|\mathbf{i}|}$ will generally be the best value to make sure that this term is bounded. Indeed, this is a requirement for the proofs of multiple results in the PAC-Bayesian theory. Thus, we will simply always choose $t=n-\operatorname{|\mathbf{i}|}$. With this choice made, we denote :
$\operatorname{\mathcal{E}}_{\Delta}(\operatorname{\mathbf{i}},\sigma)=\operatorname*{\mathbb{E}}_{T_{\operatorname{\mathbf{i}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{i}|}}}\operatorname*{\mathbb{E}}_{T_{\operatorname{\mathbf{i}}^{c}}\sim\operatorname{\mathcal{D}}^{n-\operatorname{|\mathbf{i}|}}}e^{(n-\operatorname{|\mathbf{i}|})\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{T_{\operatorname{\mathbf{i}}^{c}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{i}}},\sigma)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{i}}},\sigma)))}.$
We finish the proof by choosing the value of $\delta_{\operatorname{\mathbf{i}}}^{\sigma}$, which needs to be defined independently of $S$. Choose a set $I_{m}=\{\operatorname{\mathbf{i}}\in I:\operatorname{|\mathbf{i}|}=m\}$. As we have no information on which $\operatorname{\mathbf{i}}\in I_{m}$ will give the best results, we define a uniform distribution over all $\smallmatrixquantity(n\\ m)$ vectors in $I_{m}$, which gives a weight of $\smallmatrixquantity(n\\ m)^{-1}\forall\operatorname{\mathbf{i}}\in I_{m}$. Now, we generally consider multiple sizes of compression set
$I=\bigcup_{k=0}^{M}I_{k},$
so we need to define a probability distribution over each set $I_{k}$. We could simply choose $\tfrac{1}{M+1}$, but the probabilities would tend very fast to zero when we consider a large number $M$ of compression set sizes. It is a better choice, as discussed by Marchand and Sokolova, (2005) in Section 5.2, to choose :
$\zeta(m)=\frac{6}{\pi^{2}(m+1)^{2}},\quad\text{ with }\sum_{m=0}^{\infty}\zeta(m)=1.$
For any compression set $S_{\operatorname{\mathbf{i}}}$, we define a probability distribution $P_{M(\operatorname{\mathbf{i}})}$ over $M(\operatorname{\mathbf{i}})$ such that $\sum_{\sigma\in M(\operatorname{\mathbf{i}})}P_{M(\operatorname{\mathbf{i}})}(\sigma)\leq 1$.
Thus, for a $\delta\in(0,1)$, we define $\delta_{\operatorname{\mathbf{i}}}^{\sigma}=\matrixquantity(n\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})P_{M(\operatorname{\mathbf{i}})}(\sigma)\delta$ and
$\sum_{\operatorname{\mathbf{i}}\in I}\sum_{\sigma\in M(\operatorname{\mathbf{i}})}\delta_{\operatorname{\mathbf{i}}}$ $=\sum_{\operatorname{\mathbf{i}}\in I}\sum_{\sigma\in M(\operatorname{\mathbf{i}})}\matrixquantity(n\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})P_{M(\operatorname{\mathbf{i}})}(\sigma)\delta$
$=\sum_{m=1}^{M}\sum_{\operatorname{\mathbf{i}}\in I_{m}}\matrixquantity(n\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})\delta$
$=\sum_{m=1}^{M}\zeta(\operatorname{|\mathbf{i}|})\delta$
$\leq\delta.$
Thus, we have $1-\sum_{\operatorname{\mathbf{i}}\in I}\sum_{\sigma\in M(\operatorname{\mathbf{i}})}\delta_{\operatorname{\mathbf{i}}}^{\sigma}\geq 1-\delta$. We replace $\delta_{\operatorname{\mathbf{i}}}^{\sigma}$ by $\smallmatrixquantity(n\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})P_{M(\operatorname{\mathbf{i}})}(\sigma)\delta$ in Equation 9 to finish the proof.

### v3

@@ A.1.3 Regression problems
We trained decision trees and forests on the datasets, using P2L to train the models on one datapoint at a time. We trained the models until their validation loss hasn’t decreased for 10 or 20 epochs. We summarize this idea in Algorithm 2. We denote the RMSE as $\ell^{\textrm{RMSE}}(h,\boldsymbol{x},y)=\sqrt{(h(\boldsymbol{x})-y)^{2}}$ and the empirical risk on the dataset
@@
Input : $T$, the number of look-ahead iterations to perform before stopping.
Initialize : $S_{\operatorname{\mathbf{i}}}\leftarrow\emptyset$.
Initialize : $h_{\operatorname{\mathbf{i}}}\leftarrow h_{0}$.
Initialize : $\operatorname{\mathcal{L}}_{\textrm{BEST}}\leftarrow\infty.$
Initialize : $\texttt{COUNTER}\leftarrow 0$.
Initialize : $(\overline{\boldsymbol{x}},\overline{y})\leftarrow\operatorname*{arg\,max}_{(\boldsymbol{x},y)\in S}\ell^{\textrm{RMSE}}(h_{0},\boldsymbol{x},y)$
@@
  • Maximum depth of the trees : $\{5,10\}$
  • Minimum samples to split : $\{2,3,4\}$
  • Minimum samples to create a leaf : $\{1,2,3\}$
  • Cost-Complexity pruning parameter : $\{0.0,0.05,0.1,0.2,0.5,1,2\}$
  • Number of epochs before early stopping: $\{10,20\}$
@@
For the decision forests, we choose the number of estimators in $\{50,100\}$. For the baselines, we train the same model with the same hyperparameters on the whole dataset. We present the results for the forests in Table 4 and Table 12.
@@
Table 12: Results for the decision forests trained using P2L. We report the RMSE achieved by the models with the smallest validation loss. The ratio $\operatorname{|\mathbf{i}|}/n$ is presented in percents (%).
@@
Dataset Train loss Validation loss Test loss $\mathrm{kl}$ bound Linear bound $\operatorname{|\mathbf{i}|}/n$ Baseline test loss $\ell^{\max}$
Powerplant 4.15$\pm$0.17 4.47$\pm$0.21 4.51$\pm$0.21 18.66$\pm$1.71 25.74$\pm$2.06 2.13$\pm$0.53 3.59$\pm$0.13 90.6
Infrared 0.21$\pm$0.01 0.23$\pm$0.02 0.26$\pm$0.01 1.19$\pm$0.06 1.48$\pm$0.06 3.60$\pm$0.46 0.23$\pm$0.01 4.26
Airfoil 1.78$\pm$0.21 2.28$\pm$0.29 2.37$\pm$0.17 21.91$\pm$1.03 25.41$\pm$1.00 14.61$\pm$2.03 2.10$\pm$0.15 45.13
Parkinson 3.48$\pm$0.44 3.69$\pm$0.36 3.83$\pm$0.45 17.23$\pm$0.55 19.24$\pm$0.85 7.13$\pm$1.22 2.23$\pm$0.16 45.13
Concrete 4.32$\pm$0.49 5.54$\pm$0.71 5.43$\pm$0.57 45.58$\pm$4.43 51.71$\pm$3.69 14.87$\pm$4.19 4.70$\pm$0.36 90.63
The results for the decision trees can be found in Table 5 and Table 13. Using only P2L to train the trees leads to underfitted trees, as the model is not complex enough to use only a few datapoints to train a complete model.
Table 13: Results for the decision trees trained using P2L. We report the RMSE achieved by the models with the smallest validation loss. The ratio $\operatorname{|\mathbf{i}|}/n$ is presented in percents (%).
Dataset Train loss Validation loss Test loss $\mathrm{kl}$ bound Linear bound $\operatorname{|\mathbf{i}|}/n$ Baseline test loss $\ell^{\max}$
Powerplant 8.85$\pm$0.98 9.02$\pm$0.84 9.17$\pm$1.01 24.74$\pm$1.42 29.19$\pm$1.68 1.86$\pm$0.51 4.07$\pm$0.13 90.6
Infrared 0.24$\pm$0.02 0.27$\pm$0.03 0.31$\pm$0.04 1.76$\pm$0.05 2.08$\pm$0.04 8.39$\pm$0.48 0.27$\pm$0.03 4.26
Airfoil 6.54$\pm$0.89 6.55$\pm$0.54 6.41$\pm$0.44 28.94$\pm$0.71 30.73$\pm$0.65 15.70$\pm$2.83 3.01$\pm$0.19 45.13
Parkinson 12.19$\pm$1.82 12.20$\pm$1.73 12.09$\pm$2.01 21.91$\pm$0.48 22.18$\pm$0.55 2.47$\pm$1.31 3.20$\pm$0.15 41.37
Concrete 10.24$\pm$1.13 10.93$\pm$1.43 10.69$\pm$0.98 59.40$\pm$2.03 62.39$\pm$1.84 19.83$\pm$3.38 6.22$\pm$0.91 90.63
When computing the generalization bounds, we need to assume that the loss is bounded (for the kl bound) or sub-gaussian (for the linear bound).
Note that the RMSE is not bounded. However, the regression trees cannot predict a value greater (respectively lower) than the highest (respectively lowest) target value found in the training dataset. To compute the $\mathrm{kl}$ bound, we work under the assumption that the target space $\operatorname{\mathcal{Y}}$ is such that
${\mathcal{Y}}=\left[y_{-}-\tfrac{1}{10}\quantity(y_{+}-y_{-}),y_{+}-\tfrac{1}{10}\quantity(y_{+}-y_{-})\right]\mbox{ with }y_{-}=\min_{(\boldsymbol{x},y)\in S}y\mbox{ and }y_{+}=\max_{(\boldsymbol{x},y)\in S}y\,.$
To compute the linear bound, we assume that the distribution is $\varsigma^{2}$-sub-Gaussian with
$\varsigma=\tfrac{1}{2}\quantity(y_{+}-y_{-}).$
We report the assumed lower and greater values in Table 14.
@@
Table 14: Minimum and maximum target values used to compute the bound on regression problems.
Dataset Assumed lower Observed minimum Observed maximum Assumed upper
target value target value (in $S$) target value (in $S$) target value
@@ A.1.4 Amazon Polarity
We trained DistilBERT on the Amazon Reviews Polarity dataset. We use a subset of 10% of the real dataset, amounting to 360k datapoints. 180k are used for pretraining the model, 144k for training and 36k for validation. We use the given test set of 400k datapoints. Using P2L, we add 32 datapoints at a time to the compression set and stop the training when the validation loss hasn’t decreased in 20 iterations. The initial model $h_{0}$ is the model pretrained on the 180k datapoints. For both P2L and the baseline, which was trained for 200 epochs or until it reached 0 errors on the training dataset, we use the following hyperparameter grid :
  • Number of pretraining epochs : $\{2,5\}$
  • Pretraining learning rate : $2\times 10^{-5}$
  • Dropout probability : $\{0.1,0.2\}$
  • Training learning rate : $\{10^{-6},10^{-7},10^{-8}\}$
@@ Appendix B THEORETICAL RESULTS FROM THE LITERATURE
$\forall\operatorname{\mathbf{i}}\in I,\forall\sigma\in M(\operatorname{\mathbf{i}})\,:$ $\ R_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma))\leq 1-\exp\quantity(\frac{-1}{n-\operatorname{|\mathbf{i}|}-\kappa}\quantity[\ln\matrixquantity(n-\operatorname{|\mathbf{i}|}\\ \kappa)+\ln\matrixquantity(n\\ \operatorname{|\mathbf{i}|})+\ln\quantity(\frac{1}{\zeta(\operatorname{|\mathbf{i}|})P_{M(\operatorname{\mathbf{i}})}(\sigma)\delta})])\,,$
with $\kappa=|\mathbf{i}^{c}|R_{S_{\operatorname{\mathbf{i}}^{c}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma))$.
@@ C.1 Proof of the main result
Before proving Theorem 3, we restate Chernoff’s bound in a way that will be useful to prove Theorem 3.
@@
[…]
@@
Chernoff’s bound states that for a random variable $X$, any $t>0$ and $\epsilon>0$, we have
$\operatorname*{\mathbb{P}}\quantity(X>\epsilon)\ \leq\ e^{-t\epsilon}\operatorname*{\mathbb{E}}e^{tX}\,.$
By choosing $\delta=e^{-t\epsilon}\operatorname*{\mathbb{E}}e^{tX}$, we have
@@
$\iff$ $\epsilon=\frac{1}{t}\quantity[\ln\operatorname*{\mathbb{E}}e^{tX}+\ln\frac{1}{\delta}]\,.$
Thus, we obtain
@@
We start by defining the set of sample-compressed predictors. Given a dataset $S\sim\operatorname{\mathcal{D}}^{n}$ and $\operatorname{\mathcal{H}}$ a predictor set, we consider the following subset of $\operatorname{\mathcal{H}}$, that contains only sample-compressed predictors :
^$\widehat{\operatorname{\mathcal{H}}}_{S}\coloneqq\left\{\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma)|\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n),\sigma\in M(\operatorname{\mathbf{i}})\right\}\subseteq\operatorname{\mathcal{H}}.$
Note that for any vector of indices $\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)$ and any message $\sigma\in M(\operatorname{\mathbf{i}})$, when given a dataset $S$, we obtain a predictor $\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma)\in\widehat{\operatorname{\mathcal{H}}}_{S}$.
For a specific pair $(\operatorname{\mathbf{i}},\sigma)$, let’s study the value of $\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{S_{\operatorname{\mathbf{i}}^{c}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(S_{\operatorname{\mathbf{i}}},\sigma)))$, a realization of a random variable of mean
@@
With $\delta_{\operatorname{\mathbf{i}}}^{\sigma}\in(0,1)$ and $t>0$, using Chernoff’s bound as stated in Lemma 12, we have
@@
$\geq 1-\delta_{\operatorname{\mathbf{i}}}^{\sigma}\,.$
Thanks to the union bound, we get a bound that is valid for all pairs $(\operatorname{\mathbf{i}},\sigma)$ simultaneously,
$\operatorname*{\mathbb{P}}_{S\sim\operatorname{\mathcal{D}}^{n}}\quantity(\forall\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n),\sigma\in M(\operatorname{\mathbf{i}}):\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{S_{\operatorname{\mathbf{i}}^{c}}}(h_{\operatorname{\mathbf{i}}}^{\sigma}),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(h_{\operatorname{\mathbf{i}}}^{\sigma}))\leq\frac{1}{t}\quantity[\ln\operatorname*{\mathbb{E}}_{T_{\operatorname{\mathbf{i}}}\sim\operatorname{\mathcal{D}}^{\operatorname{|\mathbf{i}|}}}\operatorname*{\mathbb{E}}_{T_{\operatorname{\mathbf{i}}^{c}}\sim\operatorname{\mathcal{D}}^{n-\operatorname{|\mathbf{i}|}}}e^{t\Delta\quantity(\operatorname{\widehat{\mathcal{L}}}_{T_{\operatorname{\mathbf{i}}^{c}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{i}}},\sigma)),\operatorname{\mathcal{L}}_{\operatorname{\mathcal{D}}}(\operatorname{\mathscr{R}}(T_{\operatorname{\mathbf{i}}},\sigma)))}+\ln\frac{1}{\delta_{\operatorname{\mathbf{i}}}^{\sigma}}])$
$\geq 1-\sum_{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n)}\sum_{\sigma\in M(\operatorname{\mathbf{i}})}\delta_{\operatorname{\mathbf{i}}}^{\sigma}.$ (7)
Given $\delta\in(0,1)$, we set $\delta_{\operatorname{\mathbf{i}}}^{\sigma}=\matrixquantity(n\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})P_{M(\operatorname{\mathbf{i}})}(\sigma)\delta$ and we obtain
$\sum_{\operatorname{\mathbf{i}}\in I}\sum_{\sigma\in M(\operatorname{\mathbf{i}})}\delta_{\operatorname{\mathbf{i}}}^{\sigma}$ $=\sum_{\operatorname{\mathbf{i}}\in I}\sum_{\sigma\in M(\operatorname{\mathbf{i}})}\matrixquantity(n\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})P_{M(\operatorname{\mathbf{i}})}(\sigma)\delta$
$=\sum_{m=1}^{n}\sum_{\operatorname{\mathbf{i}}\in I_{m}}\matrixquantity(n\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})\delta$
$=\sum_{m=1}^{n}\zeta(\operatorname{|\mathbf{i}|})\delta$
$\leq\delta.$
Thus, we have $1-\sum_{\operatorname{\mathbf{i}}\in I}\sum_{\sigma\in M(\operatorname{\mathbf{i}})}\delta_{\operatorname{\mathbf{i}}}^{\sigma}\geq 1-\delta$. We substitute $\delta_{\operatorname{\mathbf{i}}}^{\sigma}$ by $\smallmatrixquantity(n\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})P_{M(\operatorname{\mathbf{i}})}(\sigma)\delta$ and let $t=n-\operatorname{|\mathbf{i}|}$ in Equation (7) to finish the proof. ∎
Note that the proof of Theorem 3 relies on specific choices for the values of the variables $\delta_{\operatorname{\mathbf{i}}}^{\sigma}$ and $t$ in Equation (7). The next paragraphs discuss the rationales behind these choices.
The choice of $t=n-\operatorname{|\mathbf{i}|}$.
To turn Equation (7) into a computable bound, one needs to either compute or upper bound the following term:
@@
As the set $T$ is a realization of a $n-\operatorname{|\mathbf{i}|}$ datapoints, choosing $t=n-\operatorname{|\mathbf{i}|}$ generally ensure that this term is bounded. Indeed, this is a requirement for the proofs of multiple results in the PAC-Bayesian theory.
The choice of $\delta_{\operatorname{\mathbf{i}}}^{\sigma}={n\choose\operatorname{|\mathbf{i}|}}^{-1}\zeta(\operatorname{|\mathbf{i}|})P_{M(\operatorname{\mathbf{i}})}(\sigma)\delta$.
Importantly, the value of $\delta_{\operatorname{\mathbf{i}}}^{\sigma}$ needs to be defined independently of $S$. Thus, choosing $\delta_{\operatorname{\mathbf{i}}}^{\sigma}$ is equivalent to choosing the prior distributions $P_{\operatorname{\mathscr{P}}(n)}$ and $P_{M(\operatorname{\mathbf{i}})}$ in order to obtain
$\delta_{\operatorname{\mathbf{i}}}^{\sigma}=P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})P_{M(\operatorname{\mathbf{i}})}(\sigma)\delta.$ (8)
Consider the set $I_{m}=\{\operatorname{\mathbf{i}}\in\operatorname{\mathscr{P}}(n):\operatorname{|\mathbf{i}|}=m\}$. As we have no information on which $\operatorname{\mathbf{i}}\in I_{m}$ is likely to lead to a good compression set for the reconstruction $\operatorname{\mathscr{R}}$, we define a uniform distribution over all $\smallmatrixquantity(n\\ m)$ vectors in $I_{m}$, which gives a weight of $\smallmatrixquantity(n\\ m)^{-1}\forall\operatorname{\mathbf{i}}\in I_{m}$. Now, we want consider all possible sizes of compression set
$I=\bigcup_{k=0}^{n}I_{k},$
so we need to define a probability distribution over each set $I_{k}$. We could simply choose $\tfrac{1}{n+1}$, but the probabilities would tend very fast to zero when we consider a large number $n$ of compression set sizes. It is a better choice, as discussed by Marchand and Sokolova, (2005) in Section 5.2, to choose :
$\zeta(m)=\frac{6}{\pi^{2}(m+1)^{2}},\quad\text{ for which }\sum_{m=0}^{\infty}\zeta(m)=1.$
Then, Equation (8) is applied with $P_{\operatorname{\mathscr{P}}(n)}(\operatorname{\mathbf{i}})=\smallmatrixquantity(n\\ \operatorname{|\mathbf{i}|})^{-1}\zeta(\operatorname{|\mathbf{i}|})$. To apply the theorem, one must also provide choice of prior distributions $P_{M(\operatorname{\mathbf{i}})}$ over the messages $M(\operatorname{\mathbf{i}})$ such that $\sum_{\sigma\in M(\operatorname{\mathbf{i}})}P_{M(\operatorname{\mathbf{i}})}(\sigma)\leq 1$ for all ${\mathbf{i}}\in\operatorname{\mathscr{P}}(n)$.